OpenStudy (anonymous):

would the answer to the question 5-y=3y+25 be 9

OpenStudy (anonymous):

Y=-5

OpenStudy (anonymous):

\[5-y=3y+25\] \[5-25=3y+y\] \[-20=4y\] \[y=-\frac{20}{4}=-5\]

OpenStudy (anonymous):

it wont be..9

OpenStudy (anonymous):

yh to change 25 to - 25 as i brought it over thanks

OpenStudy (anonymous):

no worries